# Template-Guided 3D Molecular Pose Generation via Flow Matching and Differentiable Optimization

### v1

@@ B.4 Pose Optimization algorithm
  • Phase 1: Hydrogen-free refinement
@@
  • Phase 2: Hydrogen-reintroduced refinement
@@
All hyperparameters were tuned empirically (Table S2) by benchmarking on $10$ diverse molecular pairs from the training set to achieve a balance between pose accuracy, steric clash avoidance, and energetic plausibility.
Input: Ligand conformation with coordinates $x$, number of optimization steps $n_{iterations}$, learning rate $lr$
@@
Table S2: Hyperparameters for two-phase pose optimization protocol. All weights were tuned empirically. $(x\rightarrow y)$ means that the weight increase linearly over iterations from $x$ to $y$.
@@ D.1 Experimental setup
D.2 Examples
Figure S9 shows examples of poses generated with 3D alignment and docking methods.
D.3 Quality of predictions
To evaluate the quality of generated poses, we employed the PoseBusters test suite [Buttenschoen et al., 2024], which checks a range of geometric and chemical criteria to ensure physically plausible ligand conformations within the binding pocket. We observed that some of the $369$ ground-truth crystallographic query poses from AlignDockBench did not fully satisfy the PoseBusters criteria, particularly the minimum distance to the protein. As a result, we restricted our evaluation to a subset of 263 molecules for which the ground-truth poses passed all PoseBusters checks.
Figure S6 presents, for each method, the percentage of molecules achieving an RMSD below $2$Å, along with the proportion that also satisfies the PoseBusters criteria. Traditional docking methods tend to yield a higher proportion of PoseBusters-valid poses. In contrast, purely LB methods like LSalign and ROSHAMBO exhibit a substantial drop in valid poses, reflecting the importance of protein context in guiding pose generation. FMA (without PO) shows a larger drop in PoseBusters-valid poses compared to FMA-PO +, highlighting the critical role of PO and the importance of sampling multiple initial poses. Nevertheless, FMA-PO + still exhibits a $12\%$ drop, primarily due to the minimum distance-to-protein check. This suggests that the method could be further improved by incorporating additional pocket-specific constraints or assigning greater weight to the pocket score during the PO stage. Additionally, integrating protein context directly into the initial pose generation with the FMA model may enhance the physical plausibility of the generated poses.
Figure S6: Percentage of accurate poses (RMSD $<2$Å) and PoseBusters-valid (PB-valid) poses for each method, evaluated on a subset of 263 molecules whose ground-truth crystallographic poses pass the PoseBusters checks.
D.4 Redocking experiments
@@
Method # of Molecules Aligned Mean RMSD (Å ↓) % of Molecules with RMSD $<$ 2Å (↑) FMA-PO 369/369 1.86 ± 1.43 68.56 $+^{*}$ 369/369 1.52 ± 1.24 82.11 FitDock 292/369 3.27 ± 4.64 54.11 / 42.82 LSalign 368/369 2.54 ± 2.00 54.35 / 54.2 ROSHAMBO 369/369 2.87 ± 2.09 48.51 rDock 369/369 2.79 ± 2.93 60.43 Vina 368/369 1.95 ± 2.54 75.82 / 75.61
@@
D.5 Ablation study
D.5.1 PO and pocket impact
In this section, we assess the impact of PO stage and the inclusion of the pocket during PO on pose prediction performance. Table S5 summarizes the results on AlignDockBench, comparing FMA (which does not incorporate PO), and both FMA-PO and FMA-PO + with and without pocket inclusion during PO.
@@ D.5.1 PO and pocket impact
FMA, exhibits the poorest performance across all metrics, highlighting the importance of the pose refinement. Moreover, including the pocket during PO enhances the quality of the generated poses, as evidenced by lower mean RMSD values and higher percentages of accurate alignments for both FMA-PO and FMA-PO +.
Table S5: Performance comparison of FMA, FMA-PO and FMA-PO + with and without pocket included in PO, evaluated on AlignDockBench in a crossdocking scenario.
@@
Method # of Molecules Aligned Mean RMSD (Å ↓) % of Molecules with RMSD $<$ 2Å (↑) FMA 369/369 1.967 ± 1.364 64.77 FMA-PO wo pocket 369/369 1.456 ± 1.412 67.75 FMA-PO 369/369 1.855 ± 1.421 69.38 $+^{*}$ wo pocket 369/369 1.697 ± 1.390 75.61 $+^{*}$ 369/369 1.618 ± 1.327 77.78
D.5.2 Impact of the number of generated conformations
@@
We assessed the impact of the number of generated poses with FMA on final pose accuracy. Figure S8 shows the percentage of molecules with RMSD below $2$Å and $1.5$Å for the best (i.e., lowest RMSD) pose of FMA-PO and FMA-PO + as a function of the number of sampled poses prior to PO. Performance increases with the number of generated poses, with higher sampling yielding a greater proportion of low RMSD predictions.

### v2

@@ B.4 Pose Optimization algorithm
  • Phase 1: Hydrogen‐free refinement
@@
  • Phase 2: Hydrogen‐reintroduced refinement
@@
The weights of the loss terms (Table S2) were empirically tuned using a validation set of ten diverse protein–ligand complexes from the training data. For each term—shape similarity, pharmacophoric complementarity, steric clash penalty, and internal energy regularization—we explored multiple weightings. The final values were chosen based on convergence behavior, pose quality, RMSD and alignment performance across the validation set.
Input: Ligand conformation with coordinates $x$, number of optimization steps $n_{iterations}$, learning rate lrlr
@@
Table S2: Hyperparameters for two‐phase pose optimization protocol. All weights were tuned empirically. $(x\rightarrow y)$ means that the weight increase linearly over iterations from $x$ to $y$.
@@ D.1 Experimental setup
D.2 Supplementary results
D.2.1 Cumulative distribution of RMSD values
In Figure S8 we report the cumulative distribution of RMSD values on AlignDockBench for each method.
Figure S8: Cumulative distribution of RMSD values: percentage of molecules with RMSD below different thresholds
D.2.2 Redocking experiments
@@
[…]
@@
D.2.3 Comparison to HarmonicFlow
We compared FMA to the recent flow-based docking method HarmonicFlow Stärk et al. [2023] using the official settings from the authors’ repository. As shown in Table S5, FMA significantly outperforms HarmonicFlow in both mean RMSD and success rate (<2<2Å), even when reporting HarmonicFlow’s best RMSD across the 10 sampled poses. Applying PO to HarmonicFlow improves results but still underperforms FMA, highlighting the higher quality of FMA ’s pose distribution and the benefit of using the template ligand to guide docking.
Table S5: Comparison to Harmonic Flow
Method # of Molecules with successful pose generation Mean RMSD (Å ↓) % of Molecules with RMSD << 2Å (↑) FMA (Top score) 369/369 1.97 ± 1.36 64.77 FMA (Top RMSD) 369/369 1.39 ± 0.85 84.55 FMA + PO++ (Top score) 369/369 1.62 ± 1.33 77.78 FMA + PO++ (Top RMSD) 369/369 1.11 ± 0.85 91.6 Harmonic-Flow (Top RMSD) 356/369 4.47 ± 2.07 2.81/2.71 Harmonic-Flow + PO ++(Top RMSD) 356/369 3.71 ± 2.22 19.94/19.24
D.2.4 Discussion about runtime
We analyzed runtime as a function of ligand size for both FMA and PO. As shown in Table S6, PO (2.7 s per pose, run on CPU) is slower than FMA (0.84 s for 10 poses, run on GPU). In both cases, runtime increases moderately with the number of heavy atoms and rotatable bonds (Pearson $\sim$0.2), without signs of exponential growth.
Table S6: Time, * GPU
Method Average runtime (s) Pearson correlation with the number of heavy atoms Pearson correlation with the number of rotatable bonds FMA* 0.84 s (for 10 poses) 0.20 0.23 PO 2.7 s (for 1 pose) 0.19 0.23
Runtime is a critical factor for practical applications. Each molecule is processed independently, so scaling across multiple compute instances is straightforward. To illustrate this, we report in Table S7 the estimated runtime and cost of processing various dataset sizes using 100 AWS g4dn.2xlarge instances in parallel. This setup could be use within a de novo drug design pipeline, where screening around 50,000 molecules is a realistic and relevant scale. For ultra-large virtual screening campaigns (e.g., tens or hundreds of millions of molecules), further acceleration and simplification of PO would be required. For example, moving to a batched optimization strategy, rather than optimizing one molecule at a time, could significantly reduce total runtime.
Table S7: Estimated runtime (in hours) and cost (USD) using 100 AWS g4dn.2xlarge instances in parallel for different numbers of molecules.
Number of molecules 50,000 $10^{5}$ $10^{6}$
Time FMA+PO (h) 0.49 0.98 9.83
Cost FMA+PO (\$) 41 82 826
D.3 Quality of predictions
Figure S10 reports of the strain energy as defined by GenBench3D Baillif et al. [2024], for each method. FMA yields higher strain energies, due to the absence of the PO module. While unrealistic conformations are a known limitation of deep learning models [Buttenschoen et al., 2024], this was one of the key motivations for introducing our PO module. Both FMA-PO and FMA-PO+ reduce strain energy.
Table S8 reports the number of molecules exhibiting clashes with protein, per method. FMA suffers from significantly more clashes, due to the absence of pocket conditioning. FMA-PO and FMA-PO+ reduce the number of clashes, confirming the effectiveness of the PO module in improving structural plausibility.
To further evaluate the quality of generated poses, we employed the PoseBusters test suite [Buttenschoen et al., 2024], which checks a range of geometric and chemical criteria to ensure physically plausible ligand conformations within the binding pocket. We observed that some of the $369$ ground-truth crystallographic query poses from AlignDockBench did not fully satisfy the PoseBusters criteria, particularly the minimum distance to the protein. As a result, we restricted our evaluation to a subset of 263 molecules for which the ground-truth poses passed all PoseBusters checks.
Figure S11 presents, for each method, the percentage of molecules achieving an RMSD below 22Å, along with the proportion that also satisfies the PoseBusters criteria. Traditional docking methods tend to yield a higher proportion of PoseBusters-valid poses. In contrast, purely LB methods like LSalign and ROSHAMBO exhibit a substantial drop in valid poses, reflecting the importance of protein context in guiding pose generation. FMA (without PO) shows a larger drop in PoseBusters-valid poses compared to FMA-PO +, highlighting the critical role of PO and the importance of sampling multiple initial poses. Nevertheless, FMA-PO + still exhibits a $12\%$ drop, primarily due to the minimum distance-to-protein check. This suggests that the method could be further improved by incorporating additional pocket-specific constraints or assigning greater weight to the pocket score during the PO stage. Additionally, integrating protein context directly into the initial pose generation with the FMA model may enhance the physical plausibility of the generated poses.
Figure S10: Strain energies of poses across methods (log scale).
Table S8: Number of molecules with protein clashes with the protein, per method.
Method 1 clash 2 clashes $\geq$ 3 clashes
FMA 37 16 20
FMA-PO 35 9 7
FMA-PO+ 21 6 2
Vina 1 0 0
FitDock 7 5 4
Co-crystallized Ligand 0 0 0
Figure S11: Percentage of accurate poses (RMSD <2<2Å) and PoseBusters-valid (PB-valid) poses for each method, evaluated on a subset of 263 molecules whose ground-truth crystallographic poses pass the PoseBusters checks.
D.4 Ablation studies
D.4.1 FMA’s contribution
We generated 10 candidate poses per method (FMA, RDKit‑centroid, Vina, LS‑align, FitDock‑LB), applied PO to each, and selected the top‑ranked pose. Results are reported in In Table S9. The RDKit‑centroid baseline places a random conformer at the binding‑site center. FMA-PO + consistently yields the lowest RMSD among all alternatives, indicating that FMA is the best initialization strategy for PO and that its accuracy is crucial.
Table S9: PO applied to different base alignments. All methods use flexible search mode and identical PO hyperparameters.
Method + PO
# of Molecules
Aligned
Mean RMSD
(Å ↓)
% of Molecules
with RMSD << 2Å (↑)
FMA + PO++ 369/369 1.62 ± 1.33 77.78
rdkit + PO++ 369/369 4.63 ± 1.94 10.33
Vina + PO++ 368/369 1.96 ± 1.87 71.47/71.27
LS-align + PO++ 357/369 2.31 ± 2.07 63.31/61.25
Fitdock LB + PO++ 365/369 2.78 ± 2.41 52.6/52.03
D.4.2 PO’s contribution
@@
In Table S10, we evaluate the impact of the PO stage and its objectives on pose prediction. FMA, exhibits the poorest performance across all metrics, highlighting the importance of the pose refinement. Moreover, including the pocket during PO enhances the quality of the generated poses, as evidenced by lower mean RMSD values and higher percentages of accurate alignments for both FMA-PO and FMA-PO +.
Table S10: Ablation studies of PO loss terms on AlignDockBench in a crossdocking scenario.
@@
Method Mean RMSD (Å ↓) % of Molecules with RMSD << 2Å (↑) FMA 1.97 ± 1.36 64.77 FMA-PO 1.86 ± 1.42 69.38 FMA-PO w/o $\mathcal{S}_{STS}$ 1.98 ± 1.38 63.96 FMA-PO w/o $\mathcal{S}_{PTS}$ 1.86 ± 1.38 67.48 FMA-PO w/o $\mathcal{S}_{\text{pocket}}$ 1.88 ± 1.41 67.75 FMA-PO w/o $\mathcal{E}_{\text{internal}}$ 2.8 ± 0.91 13.82 FMA-PO++ 1.62 ± 1.33 77.78 FMA-PO++ w/o $\mathcal{S}_{STS}$ 1.85 ± 1.32 66.67 FMA-PO++ w/o $\mathcal{S}_{PTS}$ 1.71 ± 1.33 74.25 FMA-PO++ w/o $\mathcal{S}_{\text{pocket}}$ 1.70 ± 1.39 75.61 FMA-PO++ w/o $\mathcal{E}_{\text{internal}}$ 2.98 ± 1.26 14.09
In Table S11) we applied PO to top poses from other methods. PO consistently reduced RMSD supporting its general effectiveness as a refinement stage. PO also improves chemical plausibility and geometrical validity as demonstrated in Figure S10 and Table S8
Table S11: Impact of PO on FMA, Vina, LS‐align and FitDock LB compared to the base aligners alone.
Method # Aligned Mean RMSD (Å $\downarrow$) % RMSD << 2Å ($\uparrow$)
FMA 369/369 1.97 ± 1.36 64.77
FMA + PO 369/369 1.86 ± 1.42 69.38
Vina 368/369 3.39 ± 2.81 47.28 / 47.15
Vina + PO 368/369 3.24 ± 2.85 52.17 / 52.03
LS‐align 368/369 2.54 ± 2.00 54.35 / 54.20
LS-align + PO 357/369 2.47 ± 2.15 59.66/57.72
FitDock LB 366/369 2.93 ± 2.36 49.73/ 49.32
Fitdock LB + PO 366/369 2.75 ± 2.38 52.88/52.3
D.4.3 Vina contribution to $\mathcal{S}_{\text{score}}$
We evaluated the influence of the $S_{\text{Vina}}$ component in the total score $\mathcal{S}_{\text{score}}$ by systematically ablating it across three key configurations: FMA, FMA-PO(i.e., scoring of FMA poses and PO applied to the top-ranked pose), and FMA-PO++(i.e., PO applied to all poses, and scoring of optimized poses). The results are summarized in Table S12. These results indicate that Vina scoring provides limited benefit when applied to unoptimized poses (FMA or FMA-PO), but can help select better candidates among refined poses (FMA-PO++). Our method still outperforms all baselines without using the Vina score in $\mathcal{S}_{\text{score}}$.
Table S12: Vina contribution to $\mathcal{S}_{\text{score}}$
Method
Vina term
in $\mathcal{S}_{\text{score}}$
Mean RMSD
(Å ↓)
% of Molecules
with RMSD << 2Å (↑)
FMA Yes 1.97 ± 1.36 64.77
FMA No 1.96 ± 1.35 64.77
FMA-PO Yes 1.86 ± 1.42 69.38
FMA-PO No 1.81 ± 1.40 69.38
FMA-PO++ Yes 1.62 ± 1.33 77.78
FMA-PO++ No 1.68 ± 1.38 75.61
D.4.4 Impact of the number of generated conformations
@@
We assessed the impact of the number of generated poses with FMA on final pose accuracy. Figure S12 shows the percentage of molecules with RMSD below 22Å and 1.51.5Å for the best (i.e., lowest RMSD) pose of FMA-PO and FMA-PO + as a function of the number of sampled poses prior to PO. Performance increases with the number of generated poses, with higher sampling yielding a greater proportion of low RMSD predictions.